Name the different factors that affect the period of mass oscillating on a spring. And why?


1
Expert's answer
2021-05-23T18:54:56-0400

There is the following formula for the period of mass oscillating on a spring:


"T=2\pi\sqrt{\dfrac{m}{k}},"

here, "m" is the mass and "k" is the spring constant.

As we can see, there are two factors that affect the period of mass oscillating on the spring: the mass and the spring constant. If the mass increases the period also increases (because its hard to move greater mass, so it takes longer time to complete one cycle). If the spring constant increases the period decreases (because, according to Hooke's law, greater spring constant gives us more force for the same amount of stretch, so the mass can moves quickly and it takes shorter time to complete one cycle).